In a recent benefit-cost analysis of a proposed regulation that generated positive net benefits in the present but negative net benefits in the future, the government used a 7% social discount rate. If it had instead used a 8% social discount rate, the present value of the policys net benefits would have been : present value of the policys net benefits would have been ; if it had used a 6% social discount rate, the A) Lower; lower B) Lower; higher C) Higher; lower D) Higher; higher Consider a market with the following demand and supply functions The market equilibrium price is A) 1.5 B) 2.5 C) 6 D) 10 E) 12
Answer #1

1. Ans: Lower ; higher

Explanation:

Present value of a cash flow and discount rate are inversely related. So, if discount rate is greater than 7% (i.e., 8%), it will lower the present value of net benefit and if discount rate is less than 7% (i.e., 6%), the present value of net benefit will be higher.

Thus, option [B] is the correct answer.

2. Ans: 1.5

Explanation:

For equilibrium

Demand = Supply

3 - 0.25Q = 0.25Q

0.5Q = 3

Q = 3 / 0.5 = 6

P = 3 - 0.25Q = 3 - 0.25(6) = 1.5 (equilibrium price)